solve the following homogeneous equations

  1. dy/dx=(2x-7y)/(3y-8x)
  2. (2xy+3y2)dx-(2xy+x2)dy=0
  3. xsin(y/x)(ydx+xdy)+ycos(y/x)(xdy-ydx)=0
1
Expert's answer
2021-07-19T07:45:04-0400

Solution

1."\frac{dy}{dx}=\frac{2x-7y}{3y-8x}"

Take

y=vx

"\frac{dy}{dx}=v+x\frac{dv}{dx}"

v+x"\frac{dv}{dx}=\frac{2x-7vx}{8vx-8x}"

Divide the R.H.S with x and seperate by variables

"\frac{8v-8}{v-8v^2+2}dv=\frac1xdx"

Integrate both sides respectively,we have;

"\frac{-13ln(8v^2-v-2)+3\sqrt{65}(ln(16v+\sqrt{65}-1)-ln(16-\sqrt{65}-1)}{26}=ln(x)+C"

But v="\frac yx" ,replace in the equation;

"\frac{-ln(8\frac{y^2}{x^2}-\frac yx-2)+3\sqrt{65}(ln(16\frac yx+\sqrt{65}-1)-ln(16\frac yx-\sqrt{65}-1)}{26}=ln(x)+C"

2.

(2xy+3y2)dx-(2xy+x2)dy=0

Rewrite as follows,

"\frac{2xy+3y^2}{2xy+x^2}=\frac{dy}{dx}"

Take y=vx

"\frac{dy}{dx}=v+x\frac{dv}{dx}"

Replace

"\frac{2vx^2+3v^2x^2}{2vx^2+x^2}=v+x\frac{dv}{dx}"

Divide by x2 and put like terms together.

"\frac{2v+1}{v^2+v}dv=\frac 1xdx"

Integrate both sides;

ln(v2+v)=ln(x)+C

Replace back v="\frac yx"

ln("\frac{y^2}{x^2}" +"\frac yx")=ln(x)+C

3.

Rewrite as follows;

(chain(y/X)-y2cos(y/x)dx+(x2sin(y/x)-xycos(y/x)dy=0

"\frac{dy}{dx}" ="\frac{y^2cos(\frac yx)-xysin(\frac yx)}{x^2sin(\frac yx)+xycos(\frac yx)}"

Take

y=vx and "\frac {dy}{dx}=v+x\frac {dv}{dx}"

Replace in the equation and divide x2

"v+x\frac{dv}{dx}=\frac{v^2cosv-vsinv}{sin v+vcosv}"

Seperate variables;

"\frac{sinv +vcosv}{-2vsinv}dv=\frac 1xdx"

Integrate both sides using suitable methods,we have;

"\frac{-ln(sin v)-ln( v)}{2}=ln(x)+C"

Replace back v="\frac yx"


"\frac{-ln(sin(\frac yx)-ln(\frac yx)}{2}=ln(x )+C"